Circular No. 5907

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 P. Zhao, P. Callanan, M. Garcia, and J. McClintock, Smithsonian As-

trophysical Observatory, report: “The x-ray nova GRO J0422+32, which is currently in outburst (IAUC 5901), continues to brighten optically. CCD photometry obtained with the F. L. Whipple Observatory’s 1.2-m telescope gives the following magnitudes (± 0.02): Dec. 10.35 UT, R = 14.89; 10.47, 14.81; 13.12, R = 14.92, V = 15.21; 14.09, 14.90, 15.16; 14.37, 14.87, 15.13. The nova has brightened by > 0.5 mag during the past 6 days; it is now brighter than the reported maximum of the outburst that occurred this past August (IAUC 5842, 5844).”

R. M. Wagner and M. E. Everett, Ohio State University; C. R. Shrader, Goddard Space Flight Center, NASA; and S. G. Starrfield, Arizona State University, report: “Optical spectrograms (range 420-780 nm, resolution 1 nm) were obtained with the Multiple Mirror Telescope on Dec. 14.1-14.4 UT during the current outburst (IAUC 5901). Our spectra are similar to those obtained during the August outburst (IAUC 5842) and exhibit broad Balmer absorption lines with emission cores as well as emission lines due to He I (587.5 nm) and He II (468.6 nm) superposed on a blue continuum. The N III blend at 464-465 nm is weak or absent, in contrast to our spectra obtained during the initial 1992 August outburst (IAUC 5589). Measurements of He II (468.6 nm) give an equivalent width of 0.18 nm and a FWHM of nearly 1700 km/s.”

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Users of the ephemeris option in the Computer Service are reminded

that planetary perturbations are not considered and that the computation is therefore valid only close to the epoch 1994 May 8.0 of the orbital elements tabulated for this comet (cf. IAUC 5906). There is a separate option that provides perturbed, daily ephemerides of the nine nuclei that have been well observed.

NOVA AQUILAE 1993 Further CCD photometry by H. Mikuž, Ljubljana, Slovenia (cf. IAUC

5838): Dec. 4.72 UT, V = 13.17 ± 0.02; 12.73, 13.18 ± 0.02.

1993 December 15 Daniel W. E. Green